# LaV-CoT: Language-Aware Visual CoT with Multi-Aspect Reward Optimization for Real-World Multilingual VQA

### v2

@@ Abstract.
Vision-Language Models; Multilingual Multimodal Reasoning; Visual Chain-of-Thought; Reinforcement Tuning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Companion Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAEccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing**footnotetext: This work is a preprint.
@@ 1. Introduction
Despite recent progress in vision language models (VLMs), achieving reliable reasoning in multilingual multimodal settings remains challenging. Textual Chain-of-Thought (CoT) (Wei et al., 2022) reasoning improves interpretability and supports step-by-step inference (Wei et al., 2022). However, when applied to multilingual visual question answering (mVQA), a text-only rationale often under-exploits visual cues and weakens visual grounding. Orthogonal to the choice of CoT, current VLMs for mVQA still exhibit persistent issues: (i) language inconsistency, where outputs fluctuate across languages; (ii) visual–textual misalignment, reflecting insufficient grounding between vision-translated text and visual content; and (iii) limited multilingual visual reasoning, especially for tasks requiring complex cross-lingual, multimodal inference (Alam et al., 2024; Maaz et al., 2024; Thawakar et al., 2025; Wu et al., 2025). Such limitations constrain the deployment of VLMs in international products, which demand both robustness and interpretable reasoning.
@@
On the training side, LaV-CoT adopts a two-stage paradigm combining Supervised Fine-Tuning (SFT) (Li et al., 2022; Chen et al., 2022b; Bai et al., 2025) with Language-aware Group Relative Policy Optimization (GRPO) (Lu et al., 2023; Shao et al., 2024; Liu et al., 2025). Unlike standard optimization methods, our GRPO variant is guided by verifiable multi-aspect rewards—including language consistency reward, text segments and object count reward, final answer edit distance reward and format reward—yielding stable optimization and robust reasoning across languages and modalities.
@@ 3.3. Model Training
4. Format Reward ($R_{\mathrm{Format}}$). This default reward encourages the model to generate outputs that strictly follow the expected structural format. Let $\hat{Y}$ denote the model output, and let $T$ represent the set of required tags, such as <think></think> and <answer></answer>. The reward is defined as:
$R_{\mathrm{Format}}=\begin{cases}1,&\text{if }T\subseteq\hat{Y}\\ 0,&\text{otherwise}\end{cases}$

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Companion Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAE**footnotetext: This work is a preprint.
@@ 1. Introduction
Despite recent progress in vision language models (VLMs), achieving reliable reasoning in multilingual multimodal settings remains challenging. Textual Chain-of-Thought (CoT) (Wei et al., 2022) reasoning improves interpretability and supports step-by-step inference (Wei et al., 2022). However, when applied to multilingual visual question answering (mVQA), a text-only rationale often under-exploits visual cues and weakens visual grounding. Orthogonal to the choice of CoT, current VLMs for mVQA still exhibit persistent issues: i) Limited Interpretability. Direct answers produced by recent VLMs (Bai et al., 2025; Wang et al., 2025) remain as black box, hindering model transparency and interpretability. (ii) Visual–textual misalignment. Textual CoT approaches (Wei et al., 2022; Xu et al., 2025) introduce intermediate natural language reasoning steps to enhance performance. However, these text-only rationales often under-exploit visual cues, leading to weak visual grounding. (iii) Language inconsistency. Recent Visual CoT methods (Shao et al., 2024a; Zhao et al., 2025) incorporate visual cues to enable multimodal reasoning, but still suffer from inconsistency across languages. Such limitations constrain the deployment of VLMs in international products, which demand both robustness and interpretable reasoning.
@@
On the training side, LaV-CoT adopts a two-stage paradigm combining Supervised Fine-Tuning (SFT) (Li et al., 2022; Chen et al., 2022b; Bai et al., 2025) with Language-aware Group Relative Policy Optimization (GRPO) (Lu et al., 2023; Shao et al., 2024b; Liu et al., 2025). Unlike standard optimization methods, our GRPO variant is guided by verifiable multi-aspect rewards—including language consistency reward, text segments and object count reward, final answer edit distance reward and format reward—yielding stable optimization and robust reasoning across languages and modalities.
@@ 3.3. Model Training
4. Format Reward ($R_{\mathrm{Format}}$). This reward encourages the model to generate outputs that adhere to the prescribed structural format. Let $\hat{Y}$ denote the model output, and let $T=\{(t_{i},w_{i})\}$ be the set of required tag pairs $t_{i}$ with weights $w_{i}\in[0,1]$ such that $\sum_{i}w_{i}=1$, e.g., <think></think> and <answer></answer>. The format reward is computed as a weighted sum over all tags:
[…]
, where the indicator function is defined as:
[…]